OpenStudy (anonymous):

a person walks first at a constant speed of 4.90 m/s along a straight line from point A to point B and then back along the line from B to A at a constant speed of 2.85 m/s 1.what is her average speed over the trip? 2.what is her average velocity over the entire trip?

OpenStudy (anonymous):

do you think 1 is 5.7?

OpenStudy (anonymous):

o wait jk its 7.75

OpenStudy (chaise):

If you add up her two speeds and divide by two you get her average speed of 3.875m/s. Her velocity is 0 for the reason that once she comes back to point A she hasn't moved and therefore her velocity is 0 (velocity = displacement/time)

OpenStudy (anonymous):

it says the average is wrong

OpenStudy (anonymous):

velocity is right

OpenStudy (chaise):

I'm sorry; can't help you then. Got no idea what I'm doing wrong; goodluck.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

r t = d, t=d/r, r = d/t Let d be one m, a convenient number since the distance is not defined in the problem statement. The following expression divides the total distance walked by the total time consumed for the round trip walking.\[\text{AverageSpeed}=\frac{2}{\frac{1}{4.90}+\frac{1}{2.85}}=3.60\text{ }m/s \]

OpenStudy (anonymous):

thats wrong too lol

OpenStudy (anonymous):

@ riptidej Would you be so kind as to provide the "right" answers here and justify them. Thanking you in advance.

OpenStudy (anonymous):

it doesnt tell me what is right or wrong lol

OpenStudy (chaise):

Then how do you know if they are right or wrong?

OpenStudy (anonymous):

riptidej, riptidej, Don't bailout on us now !